# Average Cost Optimality of Partially Observed MDPs: Contraction of Non-Linear Filters, Optimal Solutions and Approximations

### v2

@@ Abstract.
The average cost optimality is known to be a challenging problem for partially observable stochastic control, with few results available beyond the finite state, action, and measurement setup, for which somewhat restrictive conditions are available. In this paper, we present explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. In particular, we present a new contraction based analysis, which is new to the literature to our knowledge, building on recent regularity results for non-linear filters. Beyond establishing existence, we also present several implications of our analysis that are new to the literature: (i) robustness to incorrect priors (ii) near optimality of policies based on quantized approximations, (iii) near optimality of policies with finite memory, and (iv) convergence in Q-learning. In addition to our main theorem, each of these represents a novel contribution for average cost criteria.
@@ 1. Introduction
We study optimal control for partially observable Markov decision processes (PODMPs), under the average cost criterion. Let $\mathbb{X}$ denote a standard Borel space which is the state space of a partially observed controlled Markov process. Let $\mathbb{B}(\mathbb{X})$ be its Borel $\sigma$-field. Let $\mathbb{C}_{b}(\mathbb{X})$ be the set of all continuous, bounded functions on $\mathbb{X}$. Here and throughout the paper $\mathbb{Z}_{+}$ denotes the set of non-negative integers and $\mathbb{N}$ denotes the set of positive integers. Let $\mathbb{Y}$ be a standard Borel space denoting the observation space of the model, and let the state be observed through an observation channel $Q$. The observation channel, $Q$, is defined as a stochastic kernel (regular conditional probability) from $\mathbb{X}$ to $\mathbb{Y}$, such that $Q(\,\cdot\,|x)$ is a probability measure on the power set $P(\mathbb{Y})$ of $\mathbb{Y}$ for every $x\in\mathbb{X}$, and $Q(A|\,\cdot\,):\mathbb{X}\to[0,1]$ is a Borel measurable function for every $A\in P(\mathbb{Y})$. A decision maker (DM) is located at the output of the channel $Q$, and hence it only sees the observations $\{Y_{t},\,t\in\mathbb{Z}_{\geq 0}\}$ and chooses its actions from $\mathbb{U}$, the action space which is a compact set. An admissible policy $\gamma$ is a sequence of control functions $\{\gamma_{t},\,t\in\mathbb{Z}_{\geq 0}\}$ such that $\gamma_{t}$ is measurable with respect to the $\sigma$-algebra generated by the information variables $I_{t}=\{Y_{[0,t]},U_{[0,t-1]}\},\quad t\in\mathbb{N},\quad I_{0}=\{Y_{0}\},$ where
@@
    In the field of Partially Observable Markov Decision Processes (POMDPs), the average cost optimality equation (ACOE) typically has few findings in the literature, especially in scenarios beyond those with finite states, actions, and measurements, often restricted by stringent conditions, as we discuss in detail in this paper. This paper introduces a novel result concerning the existence of a solution for the ACOE, by utilizing a Wasserstein regularity result (Theorem 2.2), as detailed in Theorem 1.1. Under Assumption 1, we show the existence of a solution for the ACOE. We provide a detailed comparison with the existing results in the literature and highlight the explicit nature of our conditions.
@@
    Subsequently, the paper then presents several applications and implications of the existence of a solution to ACOE for POMDPs. These represent new contributions in this area of study to the best of our knowledge:
@@
      Subsection 3.2.1 shows the near-optimality of quantized approximation policies for average cost criteria.
@@
      Subsection 3.2.2 examines how, under conditions of filter stability, optimal policies derived from a finite window of measurements and actions for the discounted cost criteria are near-optimal solutions for the average cost criteria in POMDPs.
@@ 1.1. Notation and preliminaries
It is known that any POMDP can be reduced to a completely observable Markov process [1], [2], whose states are the posterior state distributions or ’beliefs‘ of the observer; that is, the state at time $n$ is
@@
We call this equivalent process the filter process. We denote by $\mathcal{Z}:={\mathcal{P}}(\mathbb{X})$ the set of probability measures on $(\mathbb{X},\mathbb{B}(\mathbb{X}))$ under the weak topology, where, under this topology $\mathcal{Z}$ is also a standard Borel space. The filter process has state space $\mathcal{Z}$ and action space $\mathbb{U}$. Let $\mathbb{P}(\mathcal{Z})$ denote the probability measures on $\mathcal{Z}$, equipped with the weak topology.
@@
This way, we obtain a completely observable Markov decision process from the POMDP, with the components $(\mathcal{Z},\mathbb{U},\tilde{c},\eta)$. For the filter process, the information variables is defined as
$\tilde{I}_{t}=\left\{Z_{[0,t]},U_{[0,t-1]}\right\},\quad t\in\mathbb{N},\quad\tilde{I}_{0}=\left\{Z_{0}\right\}$
Let $\{\mu_{n},\,n\in\mathbb{N}\}$ be a sequence in $\mathcal{P}(\mathbb{X})$. The sequence $\{\mu_{n}\}$ is said to converge to $\mu\in\mathcal{P}(\mathbb{X})$ weakly if
@@ 1.2. Average Cost Optimality Equation
Average cost optimality equation (ACOE) plays a crucial role for the analysis and the existence results of MDPs under the infinite horizon average cost optimality criteria. The triplet $(h,\rho^{*},\gamma^{*})$, where $h,\gamma:\mathcal{Z}\to\mathbb{R}$ are measurable functions and $*\rho*\in\mathbb{R}$ is a constant forms the ACOE if
@@
$\sup_{\gamma\in\Gamma}\lim_{t\to\infty}\frac{E_{z}^{\gamma}[h(Z_{t})]}{t}=0,\quad\forall z\in\mathcal{Z}$
then $\gamma^{*}$ is an optimal policy for the POMDP under the infinite horizon average cost optimality criteria, and
@@
We will refer to the function $h$, the relative value function, for the rest of the paper. Note that there may not be a unique relative value function $h$ that satisfies the ACOE, however, any $h$ that satisfies the ACOE can be used for optimality analysis.
@@ 1.3. Statement of the Main Result
    The cost function $c$ is bounded and continuous.
@@
    $K_{2}:=\frac{\alpha D(3-2\delta(Q))}{2}<1.$
@@ 1.4. Examples
Let $\mathbb{X}=[0,2]$, $\mathbb{Y}=\{0,1\}$, $\mathbb{U}=[0,12]$, let the transition kernel $\mathcal{T}(.\mid x,u)=U(0,min(2,1+((x+u)/7)))$, where $U$ stands for uniform distribution, let measurement kernel $Q(x)=\lfloor x\rfloor$ and let cost function $c(x,u)=x+u$.
@@
$\left\|T(\cdot|x,u)-T(\cdot|x^{\prime},u)\right\|_{TV}$
$\leq\left\|U(0,1+(x^{\prime}+u)/7)-U(0,1+(x+u)/7)\right\|_{TV}$
@@
We have $D=supd(x,x^{\prime})=2$, and by choosing $\alpha=\frac{2}{7}$ and $K_{1}=1$, we can set $K_{2}$ to $6/7$. This choice ensures that we satisfy the conditions outlined in Assumption 1. Consequently, we can directly apply Theorem 1.1 to this specific example.
@@
Here, $\varphi(\cdot)$ is the probability density function of the standard normal distribution and Φ()normal-$\Phi(\cdot)$ is its cumulative distribution function.
@@ 1.5. Literature Review and Comparison
In the following we present a detailed literature review and also make a comparison with our results. We will in particular highlight that many of the results in the literature are not easy to verify, which also serves to demonstrate that the problem is a challenging one.
References [7, 8, 9, 10] study the average-cost control problem under the assumption that the state space is finite; they provide reachability type conditions for the belief kernels. Reference [10] also provides a detailed literature review. An additional line of argument, via simulation and coupling, was introduced in [11, 12, 13, 14] considers the finite model setup and [13] considers the case with finite-dimensional real-valued state spaces under several technical assumptions on the controlled state process via a coupling argument. Under certain strong continuity requirements of the transition kernel, [15] extends the positivity condition to Polish spaces. We also note that, via the weak Feller property of non-linear filters, the convex analytic methods can also be utilized [16, Theorem 1.2], though the dependence on the initial condition is a limitation.
Borkar and Budhiraja’s [13] assumptions and results can be explained as follows. They consider $\mathbb{X}$, $\mathbb{Y}$, and $\mathbb{U}$ as Polish spaces, where $\mathbb{X}$ is a finite-dimensional Euclidean space (not necessarily compact, unlike ours) and $\mathbb{U}$ is a compact space. In their setup, unlike ours, the observation measurement depends on the previous time’s state and action. For $(x,u)\in\mathbb{X}\times\mathbb{U}$, the transition probability measure $p(x,u,dz,dy)\in\mathcal{P}(\mathbb{X}\times\mathbb{Y})$ is defined. [13] assumes that $p$ is continuous in a strong sense: Let $\lambda$ denote the Lebesgue measure on $\mathbb{X}$ and let $\eta\in\mathcal{P}(\mathbb{Y})$, assume a density function $\varphi(x,u,z,y)$ on $\mathbb{X}\times\mathbb{U}\times\mathbb{X}\times\mathbb{Y}$ such that $p(x,u,dz,dy)=\varphi(x,u,z,y)\lambda(dz)\eta(dy)$, with $\varphi(\cdot)>0$. It is assumed that $\varphi(x,u,z,y)$ is continuous. Note that by Scheff’e’s lemma, this implies that the transition kernel as well as the measurement kernel are continuous in total variation. The update rule is then given as
@@
Compared to this result, we present complementary conditions which are explicit and testable. Additionally, for the compact setup, we do not have continuity assumptions the measurement kernels and the condition (8) is not needed. Our paper looks to be the first one where contraction properties of filter kernels is utilized.
In [7], Platzman addresses the finite-state average-cost control problem with finite state, observation, and action spaces, under very restrictive reachability, subrectangularity, and detectability conditions.
Runggaldier and Stettner in [9] consider $X$ and $Y$ as finite and $U$ as compact. They prove that under the following positivity condition [9], the Average Cost Optimal Equation (ACOE) has a bounded solution:
@@
$\inf_{i,j\in\boldsymbol{X}}\inf_{u,u^{\prime}\in\boldsymbol{U}}\inf_{\left\{C\in\mathbb{B}(\mathbb{X}):[T(C|i,u)]>0\right\}}\frac{T\left(C|j,u^{\prime}\right)}{T(C|i,u)}>0.$
@@
Borkar [11] employed a coupling argument for the finite case. In order for the Average Cost Optimal Equation (ACOE) to possess a solution, the following is assumed.
@@
Hsu, Chuang, and Arapostathis [10] consider a finite $\mathbb{X}$ and $\mathbb{Y}$ and a compact $\mathbb{U}$. They provide two different sets of assumptions under which the Average Cost Optimal Equation (ACOE) has a solution.
@@
$\mathcal{Z}_{\epsilon}=\{\mu\in\mathcal{Z}:\mu(x)>\epsilon\quad\forall x\in X\}$. There exist constants $\epsilon>0$, $k_{0}\in\mathbb{N}$, and $\alpha<1$ such that if $z_{*}(\beta)\in\arg\min_{z\in\mathcal{Z}}J^{*}_{\beta}(z)$, then for each $\beta\in[\alpha,1)$ we have
@@
$\mathbb{P}(X_{k}=j|X_{0}=i,U_{t-1},Y_{t},1\leq t\leq k)\geq\Delta\text{ for all }i,j\in X.$
@@
There exists $k\geq 1$, and $\Delta>0$ such that, for all $y^{k}\in\mathbb{Y}^{k}$ and $u^{k}\in\mathbb{U}^{k}$,
$P_{ij}\left(y^{k}|u^{k}\right)\geq\Delta\sum_{\ell\in\boldsymbol{X}}P_{\ell j}\left(y^{k}|u^{k}\right)\quad\forall i,j\in\boldsymbol{X}$
@@
$\lambda\left(u_{1},u_{2},\mu,\nu\right):=\inf_{\{C:T(C|\mu,u_{1})>0\}}\frac{T(C|\nu,u_{2})}{T(C|\mu,u_{1})}.$
@@
$\lambda\left(u^{1},u_{n}^{2},\mu,\nu_{n}\right)\rightarrow 1$ and $\lambda\left(u_{n}^{2},u^{1},\nu_{n},\mu\right)\rightarrow 1$ when $u_{n}^{2}\rightarrow u^{1}$ and $\nu_{n}\Rightarrow\mu$. Similarly, assume that both $\lambda\left(\nu_{n},\mu\right)$ and $\lambda\left(\mu,\nu_{n}\right)$ converge to 1 as $\nu_{n}\Rightarrow\mu$.

### v3

@@ Abstract.
The average cost optimality is known to be a challenging problem for partially observable stochastic control, with few results available beyond the finite state, action, and measurement setup, for which somewhat restrictive conditions are available. In this paper, we present explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. In particular, we present a new contraction based analysis, which is new to the literature to our knowledge, building on recent regularity results for non-linear filters. Beyond establishing existence, we also present several implications of our analysis that are new to the literature: (i) robustness to incorrect priors, (ii) near optimality of policies based on quantized approximations, (iii) near optimality of policies with finite memory, and (iv) convergence in Q-learning. In addition to our main theorem, each of these represents a novel contribution for average cost criteria.
@@ 1. Introduction
We study optimal control for partially observable Markov decision processes (PODMPs) under the average cost criterion. Let $\mathbb{X}$ denote a standard Borel space, which is the state space of a partially observed controlled Markov process. Let $\mathbb{B}(\mathbb{X})$ be its Borel $\sigma$-field. Let $\mathbb{C}_{b}(\mathbb{X})$ be the set of all continuous, bounded functions on $\mathbb{X}$. Here and throughout the paper $\mathbb{Z}_{+}$ denotes the set of non-negative integers and $\mathbb{N}$ denotes the set of positive integers. Let $\mathbb{Y}$ be a standard Borel space denoting the observation space of the model, and let the state be observed through an observation channel $Q$. The observation channel, $Q$, is defined as a stochastic kernel (regular conditional probability) from $\mathbb{X}$ to $\mathbb{Y}$, such that $Q(\,\cdot\,|x)$ is a probability measure on the power set $P(\mathbb{Y})$ of $\mathbb{Y}$ for every $x\in\mathbb{X}$, and $Q(A|\,\cdot\,):\mathbb{X}\to[0,1]$ is a Borel measurable function for every $A\in P(\mathbb{Y})$. A decision maker (DM) is located at the output of the channel $Q$, and hence only sees the observations $\{Y_{t},\,t\in\mathbb{Z}_{\geq 0}\}$ and choosing its actions from $\mathbb{U}$, the action space which is a compact set. An admissible policy $\gamma$ is a sequence of control functions $\{\gamma_{t},\,t\in\mathbb{Z}_{\geq 0}\}$ such that $\gamma_{t}$ is measurable with respect to the $\sigma$-algebra generated by the information variables $I_{t}=\{Y_{[0,t]},U_{[0,t-1]}\},\quad t\in\mathbb{N},\quad I_{0}=\{Y_{0}\},$ where
@@
    In the theory of Partially Observable Markov Decision Processes (POMDPs), conditions for the existence of solutions to the average cost optimality equation (ACOE) have few findings in the literature, especially in scenarios beyond those with finite states, actions, and measurements, often restricted by stringent conditions, as we discuss in detail in the paper. This paper introduces a novel result concerning the existence of a solution for the ACOE, by utilizing a Wasserstein regularity result (Theorem 2.2), as detailed in Theorem 1.1. Under Assumption 1, we show the existence of a solution for the ACOE. We provide a detailed comparison with the existing results in the literature and highlight the explicit nature of our conditions.
@@
    Subsequently, the paper then presents several applications and implications of the existence of a solution to ACOE for POMDPs. These lead to new contributions in this area of study to the best of our knowledge:
@@
      Subsection 3.2.1 establishes near-optimality of quantized approximation policies for average cost criteria.
@@
      Subsection 3.2.2 examines how, under conditions of filter stability, optimal policies derived from a finite window of measurements and actions for the discounted cost criteria are near-optimal solutions for average cost criteria in POMDPs.
@@ 1.1. Notation and preliminaries
Belief MDP reduction for POMDPs. It is known that any POMDP can be reduced to a completely observable Markov process [1], [2], whose states are the posterior state distributions or beliefs of the observer; that is, the state at time $n$ is
@@
We call this equivalent process the filter process. We denote by $\mathcal{Z}:={\mathcal{P}}(\mathbb{X})$ the set of probability measures on $(\mathbb{X},\mathbb{B}(\mathbb{X}))$ under the weak convergence topology, where, under this topology $\mathcal{Z}$ is also a standard Borel space, that is, $\mathcal{Z}=\mathcal{P}(\mathbb{X})$ is separable and completely metrizable under the weak convergence topology. The filter process has state space $\mathcal{Z}$ and action space $\mathbb{U}$. Let $\mathcal{P}(\mathcal{Z})$ denote the probability measures on $\mathcal{Z}$, equipped with the weak convergence topology.
@@
This way, we obtain a completely observable Markov decision process from the POMDP, with the components $(\mathcal{Z},\mathbb{U},\tilde{c},\eta)$. The resulting MDP is often referred to as the belief-MDP.
Convergence notions for probability measures. Let $\{\mu_{n},\,n\in\mathbb{N}\}$ be a sequence in $\mathcal{P}(\mathbb{X})$. The sequence $\{\mu_{n}\}$ is said to converge to $\mu\in\mathcal{P}(\mathbb{X})$ weakly if
@@ 1.2. Average Cost Optimality Equation
Average cost optimality equation (ACOE) plays a crucial role for the analysis and the existence results of MDPs under the infinite horizon average cost optimality criteria. In the framework of the belief-MDP noted above, the triplet $(h,\rho^{*},\gamma^{*})$ where $h:\mathcal{Z}\to\mathbb{R}$, $\gamma^{*}:\mathcal{Z}\to\mathbb{U}$ are measurable functions and $\rho^{*}\in\mathbb{R}$ is a constant, forms the ACOE if
@@
(9) $\sup_{\gamma\in\Gamma}\lim_{t\to\infty}\frac{E_{z}^{\gamma}[h(Z_{t})]}{t}=0,\quad\forall z\in\mathcal{Z}$
then $\gamma^{*}$ is an optimal policy for the POMDP under the infinite horizon average cost optimality criterion, and
@@
We will refer to the function $h$, the relative value function, for the rest of the paper. Note that there may not be a unique relative value function $h$ that satisfies the ACOE, however, any $h$ that satisfies the ACOE and the condition (9) can be used for optimality analysis.
@@ 1.3. Statement of the Main Result
    The cost function $c$ is continuous, and thus bounded since $\mathbb{X}$ is assumed to be compact.
@@
    $K_{2}:=\frac{\alpha D(3-2\delta(Q))}{2}<1,$
    where $\delta(Q)$ is defined as in Definition 1.1.
@@ 1.4. Examples
Let $\mathbb{X}=[0,2]$, $\mathbb{Y}=\{0,1\}$, $\mathbb{U}=[0,12]$, let the transition kernel $\mathcal{T}(.\mid x,u)=\mathrm{Unif}(0,min(2,1+((x+u)/7)))$, where $\mathrm{Unif}$ stands for uniform distribution, let measurement kernel $Q(x)=\lfloor x\rfloor$ and let cost function $c(x,u)=x+u$.
@@
$\left\|\mathcal{T}(\cdot|x,u)-\mathcal{T}(\cdot|x^{\prime},u)\right\|_{TV}$
$\leq\left\|\mathrm{Unif}(0,1+(x^{\prime}+u)/7)-\mathrm{Unif}(0,1+(x+u)/7)\right\|_{TV}$
@@
We have $D=\sup d(x,x^{\prime})=2$, and by choosing $\alpha=\frac{2}{7}$ and $K_{1}=1$, we can set $K_{2}$ to $6/7$. This choice ensures that we satisfy the conditions outlined in Assumption 1. Consequently, we can directly apply Theorem 1.1 to this specific example.
@@
Here, $\varphi(\cdot)$ is the probability density function of the standard normal distribution and $\Phi(\cdot)$ is its cumulative distribution function.
@@ 1.5. Literature Review and Comparison
In the following we present a detailed literature review and also provide a comparison with our results. We will in particular highlight that many of the results in the literature are not easy to verify, which also serves to demonstrate that the problem is a challenging one.
References [7, 8, 9, 10] study the average-cost control problem under the assumption that the state space is finite; they provide reachability type conditions for the belief kernels. Reference [10] also provides a detailed literature review. An additional line of argument, via simulation and coupling, was introduced in [11, 12, 13, 14]. Under certain strong continuity requirements of the transition kernel, [15] extends the positivity condition to Polish spaces. We also note that, via the weak Feller property of non-linear filters, the convex analytic methods can also be utilized [16, Theorem 1.2], though the dependence on the initial condition is a limitation.
Borkar and Budhiraja [13] consider $\mathbb{X}$, $\mathbb{Y}$, and $\mathbb{U}$ as Polish spaces, where $\mathbb{X}$ is a finite-dimensional Euclidean space (not necessarily compact, unlike ours) and $\mathbb{U}$ is a compact space. In their setup, unlike ours, the observation measurement depends on the previous time’s state and action. For $(x,u)\in\mathbb{X}\times\mathbb{U}$, the transition probability measure $p(x,u,dz,dy)\in\mathcal{P}(\mathbb{X}\times\mathbb{Y})$ is defined. [13] assumes that $p$ is continuous in a strong sense: Let $\lambda$ denote the Lebesgue measure on $\mathbb{X}$ and let $\eta\in\mathcal{P}(\mathbb{Y})$, assume a density function $\varphi(x,u,z,y)$ on $\mathbb{X}\times\mathbb{U}\times\mathbb{X}\times\mathbb{Y}$ such that $p(x,u,dz,dy)=\varphi(x,u,z,y)\lambda(dz)\eta(dy)$, with $\varphi(\cdot)>0$. It is assumed that $\varphi(x,u,z,y)$ is continuous. Note that by Scheffé’s lemma, this implies that the transition kernel as well as the measurement kernel are continuous in total variation. The update rule is then given as
@@
Compared to this result, we present complementary conditions which are explicit and testable. Additionally, for the compact setup, we do not have continuity assumptions on the measurement kernels, and the condition (10) is not needed. Our paper looks to be the first one where contraction properties of filter kernels is utilized.
In [7], Platzman addresses the finite-state average-cost control problem with finite state, observation, and action spaces, under restrictive reachability, subrectangularity, and detectability conditions.
Runggaldier and Stettner in [9] consider $\mathbb{X}$ and $\mathbb{Y}$ as finite and $\mathbb{U}$ as compact. They prove that under the following positivity condition [9], ACOE has a bounded solution:
@@
$\inf_{i,j\in\mathbb{X}}\inf_{u,u^{\prime}\in\mathbb{U}}\inf_{\left\{C\in\mathbb{B}(\mathbb{X}):[\mathcal{T}(C|i,u)]>0\right\}}\frac{\mathcal{T}\left(C|j,u^{\prime}\right)}{T(C|i,u)}>0.$
@@
Borkar [11] employs a coupling argument under the following assumption.
@@
Hsu, Chuang, and Arapostathis [10] consider a finite $\mathbb{X}$ and $\mathbb{Y}$ and a compact $\mathbb{U}$. They provide two different sets of assumptions under which ACOE has a solution.
@@
$\mathcal{Z}_{\epsilon}=\{\mu\in\mathcal{Z}:\mu(x)>\epsilon\quad\forall x\in\mathbb{X}\}$. There exist constants $\epsilon>0$, $k_{0}\in\mathbb{N}$, and $\alpha<1$ such that if $z_{*}(\beta)\in\arg\min_{z\in\mathcal{Z}}J^{*}_{\beta}(z)$, then for each $\beta\in[\alpha,1)$ we have
@@
$\mathbb{P}(X_{k}=j|X_{0}=i,U_{t-1},Y_{t},1\leq t\leq k)\geq\Delta\text{ for all }i,j\in\mathbb{X}.$
@@
There exist $k\geq 1$, and $\Delta>0$ such that, for all $y^{k}\in\mathbb{Y}^{k}$ and $u^{k}\in\mathbb{U}^{k}$,
$P_{ij}\left(y^{k}|u^{k}\right)\geq\Delta\sum_{\ell\in\mathbb{X}}P_{\ell j}\left(y^{k}|u^{k}\right)\quad\forall i,j\in\mathbb{X},$
where $P_{ij}\left(y^{k}\mid u^{k}\right)=\mathbb{P}\left(X_{k}=j,Y_{[1,k]}=y^{k}\mid X_{0}=i,U_{[0,k-1]}=u^{k}\right)$.
@@
$\lambda\left(u_{1},u_{2},\mu,\nu\right):=\inf_{\{C:\mathcal{T}(C|\mu,u_{1})>0\}}\frac{\mathcal{T}(C|\nu,u_{2})}{T(C|\mu,u_{1})}.$
@@
$\lambda\left(u^{1},u_{n}^{2},\mu,\nu_{n}\right)\rightarrow 1$ and $\lambda\left(u_{n}^{2},u^{1},\nu_{n},\mu\right)\rightarrow 1$ when $u_{n}^{2}\rightarrow u^{1}$ and $\nu_{n}\Rightarrow\mu$. Similarly, assume that both $\lambda\left(\nu_{n},\mu\right)$ and $\lambda\left(\mu,\nu_{n}\right)$ converge to 1 as $\nu_{n}\Rightarrow\mu$ (where $\Rightarrow$ denotes weak convergence of probability measures).